### v2

# Zero-Shot Temporal Resolution Domain Adaptation for Spiking Neural Networks

Abstract
Spiking Neural Networks (SNNs) are biologically-inspired deep neural networks that efficiently extract temporal information while offering promising gains in terms of energy efficiency and latency when deployed on neuromorphic devices. However, SNN model parameters are sensitive to temporal resolution, leading to significant performance drops when the temporal resolution of target data during deployment is not the same with that of the pre-deployment source data used for training, especially when fine-tuning with new data is not possible during deployment. To address this challenge, we propose three novel domain adaptation methods for adapting neuron parameters to account for the change in time resolution without re-training on target time-resolution. The proposed methods are based on a mapping between neuron dynamics in SNNs and State Space Models (SSMs) and are applicable to general neuron models. We evaluate the proposed methods under spatio-temporal data tasks, namely the audio keyword spotting datasets SHD and MSWC as well as the neuromorphic image classification NMINST dataset. Our methods provide an alternative to - and in majority of the cases significantly outperform - the existing reference method that simply scales the time constant. Moreover, our results show that high accuracy on high temporal resolution data can be obtained by time efficient training on lower temporal resolution data and model adaptation.
keywords:
domain adaptation, temporal resolution, neuromorphic, spiking neural networks (SNNs), state-space models (SSMs).
journal: Neural Networks
[…]
1 Introduction
Domain adaptation addresses the scenarios where a model is confronted by a domain shift after deployment, while its task remains the same, such as face recognition under different lightning conditions or poses [domain_adaptation_brief_review, liang2023comprehensivesurveytesttimeadaptation]. The domain shift refers to a change in feature space between the source domain on which the model was trained, and the target domain on which the model is evaluated. The change in feature space has been studied from various aspects, particularly for visual data [Patel_2015, zubić2024state], but has rarely been explored with respect to temporal resolution shift for temporal data. Our work focuses precisely on this under-explored aspect, investigating feature discrepancies caused by differences in temporal resolution between the source and target data. While domain adaptation typically involves transferring knowledge between the source and target data by re-training parts of the model or the entire model using samples from target domain, [Patel_2015, Sun_2019_CVPR], we focus on the extreme case of no retraining on target domain.
In a conventional data acquisition set-up, where values of signals of interests are recorded at regular intervals, temporal resolution is primarily determined by the time interval between successive samples. For instance, the frame rate of a video stream or the sampling rate of an audio recording determines the temporal resolution. In the case of event-based sensing, e.g. [Gallego2022], where changes in the scene are recorded with time-stamps, temporal resolution is determined by the temporal accuracy of time-stamps.
Various practical limitations lead to data with varying temporal resolution. For instance, limited memory capacity and power constraints often require lower sampling rates of the data acquisition or storage devices [Dieter_2005, ur2016BigDataCompress]. Additionally, the quality of sensors and bandwidth limitations further restrict temporal resolution, as high-resolution sensors and real-time transmission demand more resources [chen2021distributed, park2021communication]. Consequently, changes in the temporal resolution of data has been the focus of a wide range of studies, such as increasing of temporal resolution of videos [Zuckerman2020], optimization of sensor networks’ sensing rates [Ankur2004], processing of irregularly sampled data [rubanova2019latent], optimal combination of high- and low-rate time-series data [Bae2024]. In the case of event-based cameras, effect of bandwidth and event-rate constraints on the temporal resolution have been studied [gehrig2022highresolutioneventcamerasreally, zubić2024state].
In this paper, we are interested in the scenario where the temporal resolution of the source data which the model has been trained on, and the target data which the model is to be used on is different. We investigate both the scenario of Fine-to-Coarse deployment, where the change is from high to low resolution, and Coarse-to-Fine deployment, where the change is from low to high resolution. We consider the setting where it is not possible to fine-tune the model at the edge on the data with the target temporal resolution, for instance when real-time inference is needed on the current sample from target domain. Hence, we consider temporal resolution domain adaptation with no re-training on the target dataset, see Figure 1.
Figure 1: Overview of the set-up. We investigate how models trained on data with a given source temporal resolution can be adapted for data with a different target temporal resolution, both in the Fine-to-Coarse and Coarse-to-Fine deployment directions. The diagram illustrates Coarse-to-Fine scenario.
Applications of these scenarios appear at different edge processing scenarios, where small resource-constrained devices with different sensing and computing capabilities must process their data at the edge of a wireless system with varying degrees of support from a central processing or coordination unit [chen2021distributed, park2021communication]. A possible application of Coarse-to-Fine deployment is encountered in communication constrained distributed scenarios with a central learning unit. In such scenarios, sensor data is sent to a central learning unit with reduced temporal resolution due to bandwidth or power limitations [chen2021distributed, park2021communication]. As a result, the central model is trained on coarse data and then deployed to edge devices, where higher-resolution data is available for inference. A possible application of Fine-to-Coarse deployment is encountered when the sensors have to switch to low sampling rates to reduce power consumption [Dieter_2005, Ankur2004]. This leads to having target data at edge with coarser temporal resolution than the source data used for training the deployed models at the central learning unit.
We consider the above temporal domain adaptation problems under Spiking Neural Networks (SNN) models. SNNs are biologically-inspired neural networks with stateful recurrent neurons and binary spike-based communication [MAASS19971659]. SNNs provide an attractive framework for processing data with encoded temporal information due to their ability to represent and learn spatiotemporal features [bittar2022surrogate, yik2024neurobench, Braille_SNN, eshraghian2023trainingspikingneuralnetworks, ma2025spikingneuralnetworkstemporal]. Like most machine learning algorithms, the current mainstream pipeline for SNNs relies on training on high-power computational units, mostly GPUs, and then deploying and running on low-power edge devices [10177729]. SNNs, when deployed at neuromorphic edge hardware, are especially valuable because they are designed to mimic the brain’s energy-efficient processes, offering significant reductions in power consumption and latency [9395703, caccavella2023lowpower, Braille_SNN]. SNNs are particularly effective in event-based systems, where event-based sensors, e.g. event-cameras record changes only as they occur, rather than at fixed time intervals, leading to optimized energy use and reduced data redundancy [Gallego2022].
A common practical training set-up for SNNs is the standard digital clock-based SNNs, where the temporal resolution of the data corresponds directly to the processing resolution at which the model operates. In this setup, the possible discrepancies in the temporal resolution of the source and target data constitutes a significant challenge since trained SNN model parameters are strongly dependent on the temporal resolution of source data. Hence, an SNN model that has been trained using data with a given temporal resolution may not necessarily perform well in inference on data from another sensory source even if it is recording of the same phenomena, e.g., see Section LABEL:sec:numericalResults.
Training on sequences with high temporal resolution is desired due to possible inclusion of more information and hence better performance of the model [gehrig2024low]. However this comes with a cost as difficulties arise when training SNNs on long sequences such as exploding and vanishing gradients and other instabilities due to the recurrent and saturating nature of SNNs [zucchet2024recurrentneuralnetworksvanishing]. Another limitation while training on long sequences is the significant training time as the temporal resolution dimension linearly increases the training time and computational complexity [pereznieves2022sparse]. Hence, the development of temporal adaptation methods enables training models in a fast and stable manner on low temporal resolution data, while still ensuring high performance during inference on high temporal resolution data. This setting supports energy, time and computation efficient training pipeline.
Considering both the temporal domain adaptation problem as well as challenges of training on sequences with fine temporal resolution, the main contributions of this paper are as follows:
  • We propose three novel methods for adapting SNN parameters that govern the neuron dynamics for changes in temporal resolution of the data without using any data samples with new temporal resolution.
  • The performance of the proposed methods is demonstrated using the audio datasets SHD [shddataset] and MSWC [mswcdataset] and the vision NMNIST dataset [orchard2015converting] .
  • The results show that in a wide range of scenarios, e.g. coarse-to-fine scenarios for all the above datasets, our proposed methods provide significant performance gains compared to using no temporal resolution domain adaptation method or using the existing benchmark method from the literature.
  • We show that utilizing low temporal resolution version of the data for pre-training constitutes a promising training approach for improving computational training efficiency in SNNs.
2 Related Work
The Leaky-Integrate-and-Fire (LIF) neuron model is one of the most simple and widely used neuron models in the context of SNNs. When input stimuli in form of spikes are applied to the neuron, the membrane potential increases and when a certain threshold is reached it produces an output spike. On the other hand, in the absence of input stimuli, the membrane potential decays exponentially [bittar2022surrogate], see Figure 2. The vanilla LIF neuron has only one learnable parameter, which is the membrane potential decay (the leak), and this parameter has explicit exponential dependence on the time step length of the data. Hence when dealing with change of temporal resolution, intuitive way to adapt the LIF neuron dynamic is to scale the leak parameter according to the change. In [he2020comparing] the authors explore this method of temporal domain adaptation using the simple LIF neuron. On the other hand, in [caccavella2023lowpower], the authors use Integrate-and-Fire (IF) neuron which is a special simpler case of LIF neuron without leak. Hence alternative to parameters adaptation, they employ a graded spike network and different normalization techniques for domain adaptation with promising results. However, in order to capture more complex temporal information in data, more complex neuron models such as the adaptive LIF (adLIF) neurons have shown superior performance [bittar2022surrogate, baronig2025advancingspatiotemporalprocessingspiking, adLIFNature, ma2025spikingneuralnetworkstemporal]. In the case of more complex neuron models however, the parameter might not have explicit dependence on the time-length of the data, or even any dependence at all.
In this case, it is unclear how these parameters need to be scaled to depict the change of temporal resolution, and to the best of our knowledge, this issue has not been thoroughly addressed in the existing literature.
Figure 2: Feedforward SNN model architecture consisting of four layers, each with arbitrary number of neurons. An example neuron dynamics given as a zoom-in on a neuron.
An important building block of our proposed approach is State Space Models (SSMs), which is a general framework for modelling dynamic systems. Their neural network implementations, i.e. deep SSMs, raised as an attractive alternative to the conventional deep recurrent models. It was shown that for appropriate conditioning on the state matrix dynamics, deep SSMs can handle long-range dependencies while offering stable and computationally efficient training [wang2024state]. Limited explorations of temporal resolution change have been performed using SSMs and traditional data [smith2022simplified, Table 2]. Deep SSMs have been used for event-based data in a number of works [schöne2024scalable, zubić2024state] Similar to the leak adaptation in LIF neurons, explicit correlation can be addressed between significantly simplified version of SSM and the temporal dynamics between spikes. [schöne2024scalable] explores this concept by processing incoming spikes with adapted leak based on the time-length between consecutive spikes. Alternatively, [zubić2024state] explored varying-resolution aware training of deep SSMs for vision dataset, where the training includes masking inputs outside specific frequency ranges and incorporating an additional loss function to penalize high-frequency activities.
Interesting intersection of work is emerging between SNNs and deep SSMs. One common line of work is embedding spiking dynamics in deep SSMs where at each time step the output activity of the SSM is processed by a firing threshold function [stan2023learning, bal2024rethinkingspikingneuralnetworks, karilanova2025statespacemodelinspiredmultipleinput]. Alternatively, in [du2024spiking] they combine SNN with SSMs by adding a spiking layer to the SSM. Nevertheless, the challange of drawing lines of correspondence between SNNs and linear SSMs is still present due to the non-linearity when it comes to SNNs. Our proposed methods for temporal domain adaptation for SNNs are based on this drawn parallel and the theory developed for adaptation in SSMs.
3 Problem Statement
In Section 3.1 we present a popular neuron model commonly used in the SNN literature to establish a foundation for the generalized neuron model in Section 3.2. We outline the temporal domain adaptation problem that this article focuses on in Section 3.3.
3.1 Popular spiking neuron model
The discrete-time adaptive Leaky-Integrate and Fire (adLIF) neuron is defined as [bittar2022surrogate]
$\!u[t+1]=$ $\alpha(u[t]-\theta s_{out}[t])+(1-\alpha)(\bm{W}\bm{s}_{in}[t]+\bm{V}\bm{s}_{out}[t])$
$-(1-\alpha)v[t]$ (1a)
$\!v[t+1]=$ $au[t]+\beta v[t]+bs_{out}[t]$ (1b)
$s_{out}[t]=$ $g_{\theta}(u[t])$ (1c)
where $\bm{s}_{in}\in\mathbb{Z}^{L_{b}\times 1}$ are the input spikes from the neurons in the previous layer and $\bm{W}\in\mathbb{R}^{1\times L_{b}}$ is the weight matrix associated with their connections to the neuron of observation, while $\bm{s}_{out}\in\mathbb{Z}^{L_{c}\times 1}$ are the output spikes of all the other neurons in the current layer and $\bm{V}\in\mathbb{R}^{1\times L_{c}}$ with zero diagonal is their connections to the neuron of observation. Here, $L_{c}$ and $L_{b}$ are the number of neurons in the current and the previous layer, respectively. Furthermore, $g_{\theta}$ describes the spiking behavior with firing threshold $\theta\in\mathbb{R}^{1\times 1}$; $\alpha=\exp(\Delta/\tau_{mem})$ represents the decay of the membrane potential over time such that $\Delta$ corresponds to the duration of a timestep and $\tau_{mem}$ is a membrane time constant; $\beta=\exp(\Delta/\tau_{syn})$ represents the decay of the adaptive variable with adaptation constant $\tau_{syn}$; the parameters $a$ and $b$ are without explicit dependence on the discrete time step length $\Delta$. All parameters $\alpha,\beta,a,b$ are generally learnable during training within given ranges.
Note, the weight vector $\bm{V}$ corresponds to recurrent connections between the neurons in the same layer. Recurrence represented by $\bm{V}$ is a characteristic of the network rather than the neuron.
3.2 General spiking neuron model
In this section we introduce a discrete-time general n-dimensional spiking neuron model which generalizes the two-dimensional adLIF neuron (1). This generalized model is central to the proposed methods of this paper. It is defined as
$\!\bm{v}[t+1]$ $=\bm{H}_{v}\bm{v}[t]+\bm{H}_{f}s_{out}[t]+\bm{H}_{i}\bm{W}\bm{s}_{in}[t]+\bm{H}_{r}\bm{V}\bm{s}_{out}[t]$ (2a)
$s_{out}[t]$ $=g_{\bf{\Theta}}(\bm{v}[t])=\begin{cases}1\text{ if }\bm{v}[t]\in\bf{\Theta}\\ 0\text{ otherwise}\end{cases}$ (2b)
where $\bm{v}\in\mathbb{R}^{n\times 1}$ is the state variable, while $\bm{H}_{v}\in\mathbb{R}^{n\times n}$ and $\bm{H}_{f},\bm{H}_{i},\bm{H}_{r}\in\mathbb{R}^{n\times 1}$ are matrices containing the parameters of the individual neuron. The subscripts of the $\bm{H}$ matrices, i.e. $v,f,i$, and $r$, indicate state, feedback, input and recurrence, respectively. The spiking behaviour is described by $g_{\bf{\Theta}}(\cdot)$, where the neuron spikes when the state enters into the region described by the region $\bf{\Theta}$.
We note that the adLIF neuron can be expressed as
$\!\begin{bmatrix}u[t+1]\\ v[t+1]\end{bmatrix}=$ $\begin{bmatrix}\alpha&-(1-\alpha)\\ a&\beta\end{bmatrix}\begin{bmatrix}u[t]\\ v[t]\end{bmatrix}+\begin{bmatrix}-\alpha\theta\\ b\end{bmatrix}s_{out}[t]$
$+\begin{bmatrix}1-\alpha\\ 0\end{bmatrix}\bm{W}\bm{s}_{in}[t]+\begin{bmatrix}1-\alpha\\ 0\end{bmatrix}\bm{V}\bm{s}_{out}[t].$ (3)
Thus, the adLIF neuron (1) with two state variables $\bm{v}=\begin{bmatrix}u,v\end{bmatrix}^{T}$ is a special case of the general model where
$\bm{H}_{v}$ $=\begin{bmatrix}\alpha&-(1-\alpha)\\ a&\beta\end{bmatrix},\bm{H}_{f}=\begin{bmatrix}-\alpha\theta\\ b\end{bmatrix},\bm{H}_{i}=\begin{bmatrix}1-\alpha\\ 0\end{bmatrix},\text{and}$
$\bm{H}_{r}$ $=\begin{bmatrix}1-\alpha\\ 0\end{bmatrix}.$
3.3 Problem Formulation - Temporal resolution domain adaptation for general spiking neuron model
3.3.1 Event-based data
We first provide preliminaries for the notion of temporal resolution for event-based data. Event-based data can be represented as asynchronous or synchronous. In the asynchronous representation, data is provided as a list of events with each their precise time of occurrence, while in the synchronous representation, data is provided in a time-binned format as spike sequences of a given length, each bin being populated with all events occurring during this period of time. The first type of representation is naturally adopted when data comes from event-based sensors, such as NMNIST [orchard2015converting] and the second type of data representation is typically used after post-processing of data obtained by conventional sensors, such as Spiking Heidelberg Digits (SHD) dataset [shddataset], or by post-processing of data from event-based sensors.
The standard approach for processing event-based data with SNNs is to use digital clock-based SNNs, which require the data to be in an effectively synchronous form for processing. Hence, if the data is asynchronous, it is converted into a synchronous form, such as by binning events on regular time intervals.
3.3.2 Temporal resolution
In the below, we base our exposition on synchronous data representation. Consider a continuous-time signal $\bm{s}(t)$, $0\leq t\leq t_{c}$ with a duration $t_{c}$ seconds, representing a phenomena of interest, such as a spoken keyword. We denote by $\bm{s}\in\mathbb{Z}^{N\times 1}$ an integer-valued discrete-time sequence extrapolated from $\bm{s}(t)$ where $N$ is the number of time steps. Hence, for a regular discretization of time-axis, which we focus on here, a single time step in $\bm{s}$ will represent a time duration of $\Delta_{N}=t_{c}/N$ seconds.
Suppose that we have two discrete-time data sequences associated with the same continuous signal, each with a different number of time steps. Let us refer to them as $\bm{s}_{S}\in\mathbb{Z}^{N_{S}\times 1}$ and $\bm{s}_{T}\in\mathbb{Z}^{N_{T}\times 1}$ with $N_{S}\neq N_{T}$, and thus $\Delta_{S}\neq\Delta_{T}$. Hence, $\bm{s}_{S}$ and $\bm{s}_{T}$ correspond to sequences with different time resolutions. The sequences $\bm{s}_{S}$ and $\bm{s}_{T}$ can be obtained from $\bm{s}(t)$ in various ways. One is to obtain both of the sequences directly from $\bm{s}(t)$ using different encoding parameters. Another is to obtain $\bm{s}_{S}$ directly from $\bm{s}(t)$, while obtaining $\bm{s}_{T}$ by transforming $\bm{s}_{S}$, or vice versa.
3.3.3 Model adaptation
We consider the scenario where the SNN model is pre-trained on data with time resolution $\Delta_{S}$ but we would like to use this model on data with temporal resolution $\Delta_{T}$, where $\Delta_{T}\neq\Delta_{S}$. Accordingly, we refer to training dataset as source data, and the later as target data. Similarly, we refer to $\Delta_{S}$ and $\Delta_{T}$ as temporal resolution of source and target data, respectively.
During the SNN training, two groups of parameters are learned for each neuron, parameters of the neuron dynamic, $\bm{H}_{v,(S)},\bm{H}_{f,(S)},\bm{H}_{i,(S)},\bm{H}_{r,(S)}$, and the weights of the synaptic connections between the neurons, $\bm{W},\bm{V}$. Our main goal in this article is to adjust the model parameters such that we obtain high-accuracy results on target data. We propose to achieve this goal by keeping the dynamics of the state variable $\bm{v}$ of neurons, see (2), as similar as possible under the source and target data. This means the state variable $\bm{v}$ dynamics is preserved even when the temporal resolution of the data changes.
We consider this problem under two constraints: First, we are only allowed to adapt the neuron dynamic parameters, $\bm{H}_{k}$ for $k=v,i,f,r$. Hence, SNN architecture and synaptic connections are unchanged. Second, the only known information about the target data is the ratio between the target and source temporal resolution. Hence, we do not rely on any post-deployment re-training on the target dataset, which can be costly and prone to interference. Our results in Section LABEL:sec:numericalResults show that promising performance can be obtained even under this constrained scenario, supporting the central role of neuron dynamics for capturing time-dependent behaviour.
To summarize, the main goal of this article is to find a mapping from the source to target feature space in the form of
$\bm{H}_{v,(T)},$ $\bm{H}_{i,(T)},\bm{H}_{f,(T)},\bm{H}_{r,(T)}$
$={\cal{M}}(\rho,\bm{H}_{v,(S)},\bm{H}_{i,(S)},\bm{H}_{f,(S)},\bm{H}_{r,(S)}),$ (4)
where the mapping ${\cal{M}}(.)$ represents the temporal resolution domain adaptation, and the parameter $\rho$ is given by
$\rho=\frac{\Delta_{T}}{\Delta_{S}}=\frac{N_{S}}{N_{T}}=\frac{1}{\bar{\rho}}$ (5)
defining the ratio between the temporal resolutions. Here, we have also defined $\bar{\rho}$, the reciprocal of $\rho$, for a more convenient notation in later chapters. We consider the following scenarios:
  • Coarse-to-Fine deployment: The model is pretrained on source data with coarse temporal resolution and tested on target data with fine temporal resolution, i.e., $\Delta_{S}>\Delta_{T}$, thus $\rho<1$.
  • Fine-to-Coarse deployment: The model is pretrained on source data with fine temporal resolution and tested on target data with coarse temporal resolution, i.e., $\Delta_{S}<\Delta_{T}$, thus $\rho>1$.
4 Methods
As a background for our work, in Section 4.1, we begin by expressing the SNN neuron model as a non-linear SSM model. This provides a foundation for Section 4.2, where we propose an approximate correspondence between the SNN neuron model and linear SSM model. The proposed temporal resolution domain adaptation methods, presented in Section 4.3, are based on this approximate correspondence. In Section 4.4 we present a simple benchmark method.
4.1 Preliminaries - Correspondence between SNNs and non-linear SSMs
In this section, we express the SNN neuron model as a specific case of non-linear SSMs.
A general discrete-time nonlinear SSM can be written as [OkuyamaYoshifumi2014DCS]:
$\bm{v}[t+1]$ $=\mathcal{F}(\bm{v}[t],\bm{f}[t])$ (6a)
$\bm{y}[t]$ $=\mathcal{G}(\bm{v}[t],\bm{f}[t])$ (6b)
where $\bm{v}[t]\in\mathbb{R}^{n\times 1},\bm{f}[t]\in\mathbb{R}^{r\times 1}$, $\bm{y}[t]\in\mathbb{R}^{p\times 1}$ represent the state vector, the input to the system and the output of the system, respectively, and $\mathcal{F}(\cdot)$ and $\mathcal{G}(\cdot)$ are possibly non-linear functions. Here $n$ is the dimension of the system state, $r$ is dimension of the system input, and $p$ is dimension of the system output. The general neuron model in (2) can be expressed as a non-linear SSM as in (6) using
$\!\mathcal{F}(\bm{v}[t],\bm{f}[t])$ $=\bm{H}_{v}\bm{v}[t]+\bm{H}_{f}g_{\bf{\Theta}}(\bm{v}[t])+\begin{bmatrix}\bm{H}_{i},\bm{H}_{r}\end{bmatrix}^{T}\bm{f}[t]$ (7a)
$\!\mathcal{G}(\bm{v}[t],\bm{f}[t])$ $=g_{\bf{\Theta}}(\bm{v}[t])$ (7b)
where $\bm{f}[t]=\begin{bmatrix}\bm{W}\bm{s}_{in}[t],\bm{V}\bm{s}_{out}[t]\end{bmatrix}^{T}$ is considered external input since $\bm{s}_{in}[t]$ parameterize the spikes from the previous layer and $\bm{s}_{out}[t]$ parameterize the spikes from all the other neurons in the current layer. Here, $\bm{y}$ represents the output spike of the neuron.
Since our ultimate goal is to find a correspondence between SNNs and linear SSMs, we move one step closer to a linear SSM by considering separable $\mathcal{F}(\cdot)$ and $\mathcal{G}(\cdot)$, i.e.,
$\bm{v}[t+1]$ $=\mathcal{F}_{A}(\bm{v}[t])+\mathcal{F}_{B}(\bm{f}[t])$ (8a)
$\bm{y}[t]$ $=\mathcal{G}_{C}(\bm{v}[t])+\mathcal{G}_{D}(\bm{f}[t])$ (8b)
where $\mathcal{F}_{A}(\cdot),\mathcal{F}_{B}(\cdot),\mathcal{G}_{C}(\cdot)$ and $\mathcal{G}_{D}(\cdot)$ are possibly non-linear functions. The generalized neuron (2) can be expressed as a non-linear SSM with separable functions (8) using
$\mathcal{F}_{A}(\bm{v}[t])$ $=\bm{H}_{v}\bm{v}[t]+\bm{H}_{f}g_{\bf{\Theta}}(\bm{v}[t])$ (9a)
$\mathcal{F}_{B}(\bm{f}[t])$ $=\begin{bmatrix}\bm{H}_{i},\bm{H}_{r}\end{bmatrix}^{T}\bm{f}[t]$ (9b)
$\mathcal{G}_{C}(\bm{v}[t])$ $=g_{\bf{\Theta}}(\bm{v}[t])$ (9c)
$\mathcal{G}_{D}(\bm{f}[t])$ $=\bm{0}$ (9d)
where $\bm{f}[t]=\begin{bmatrix}\bm{W}\bm{s}_{in}[t],\bm{V}\bm{s}_{out}[t]\end{bmatrix}^{T}$.
4.2 Preliminaries - Correspondence between SNNs and linear SSMs
As linear SSMs enable analytical investigations and form the basis of our adaptation methods, in this section, we propose an approximate correspondence between the general SNN neuron model in (2) and linear SSMs. Despite SNNs being strictly non-linear SSMs, this approximate correspondence can be utilized to achieve promising results for temporal domain adaptation, as demonstrated by the results in Section LABEL:sec:numericalResults.
A linear SSM in discrete-time can be written as [Gajic]:
$\bm{v}[t+1]=\bm{A}\bm{v}[t]+\bm{B}\bm{f}[t]$ (10a)
$\bm{y}[t]=\bm{C}\bm{v}[t]+\bm{D}\bm{f}[t],$ (10b)
where the vectors $\bm{v}[t],\bm{f}[t],\bm{y}[t]$ are respectively the state vector, input vector and output vector. Here, the matrix $\bm{A}\in\mathbb{R}^{n\times n}$ describes the internal behaviour of the system, while matrices $\bm{B}\in\mathbb{R}^{n\times r},\bm{C}\in\mathbb{R}^{p\times n},\bm{D}\in\mathbb{R}^{p\times r}$ represent connections between the external world and the system. These matrices represent arrays of real scalar numbers. The system is assumed to be time invariant which means the scalar numbers are constant over time.
To find an approximate correspondence between the general neuron model in (2) and (10), we first map the nonlinear function in the separable non-linear SSM (8) to the matrices describing linear relation in (10). The mapping is intuitive,
$\bm{A}\bm{v}[t]$ $\rightarrow\mathcal{F}_{A}(\bm{v}[t])$ (11a)
$\bm{B}\bm{f}[t]$ $\rightarrow\mathcal{F}_{B}(\bm{f}[t])$ (11b)
$\bm{C}\bm{v}[t]$ $\rightarrow\mathcal{G}_{C}(\bm{v}[t])$ (11c)
$\bm{D}\bm{f}[t]$ $\rightarrow\mathcal{G}_{D}(\bm{f}[t])$ (11d)
where the symbol $\rightarrow$ indicates that the left-hand side linear term is best described as part of the dynamics of the right-hand side possibly nonlinear function. Next, we associate the general neuron matrices $\bm{H}_{v},\bm{H}_{f},\bm{H}_{i},\bm{H}_{r}$ in (2), to the linear matrices $\bm{A},\bm{B},\bm{C},\bm{D}$ in (10). Using the correspondence in (9) and (11), the following mapping arises as a candidate: $\bm{A}=\begin{bmatrix}\bm{H}_{v},\bm{H}_{f}\end{bmatrix}^{T}$, and $\bm{B}=\begin{bmatrix}\bm{H}_{i},\bm{H}_{r}\end{bmatrix}^{T}$. However, we propose to make a deviation from this mapping and interpret the spike of the previous time step as external input to the neuron state. Hence, we use $\bm{A}=\bm{H}_{v}$, and $\bm{B}=\begin{bmatrix}\bm{H}_{f},\bm{H}_{i},\bm{H}_{r}\end{bmatrix}^{T}$. To visualize this correspondance, we rewrite the general spiking neuron in (2) as
$\!\bm{v}[t+1]$ $=\bm{H}_{v}\bm{v}[t]+\begin{bmatrix}\bm{H}_{f}\\ \bm{H}_{i}\\ \bm{H}_{r}\end{bmatrix}\begin{bmatrix}s_{out}[t]\\ \bm{W}\bm{s}_{in}[t]\\ \bm{V}\bm{s}_{out}[t]\end{bmatrix}$ (12a)
$s_{out}[t]$ $=g_{\bf{\Theta}}(\bm{v}[t])$ (12b)
Hence, we propose the following approximate correspondence between the general SNN neuron model and the linear SSM framework:
$\bm{v}[t+1]$ $=\bm{A}\bm{v}[t]+\bm{B}\bm{f}[t]$ (13a)
where $\bm{v}\in\mathbb{R}^{n\times 1},\bm{A}=\bm{H}_{v}\in\mathbb{R}^{n\times n},\bm{B}=\begin{bmatrix}\bm{H}_{f},\bm{H}_{i},\bm{H}_{r}\end{bmatrix}^{T}\in\mathbb{R}^{n\times r}$, and $\bm{f}[t]=\begin{bmatrix}s_{out}[t],\bm{W}\bm{s}_{in}[t],\bm{V}\bm{s}_{out}[t]\end{bmatrix}^{T}\in\mathbb{R}^{r\times 1}$ such that $\bm{f}$ combines all input spikes to the neuron including the effect of its own output. Since only the $\bm{A}$ and $\bm{B}$ matrices are needed for adapting the state variable dynamics, which is our main interest here, we leave (9c) and (9d) as they are, since these functions are associated with the output spikes $\bm{y}[t]$ and are not required for adaptation of the state variable.
4.3 Proposed Methods for Temporal Resolution Domain Adaptation
In this section, we define our proposed temporal resolution domain adaptation methods for the generalized SNN neuron model in (2). As explained in Section 3.3, the proposed methods suggest a mapping ${\cal{M}}(.)$ from the source domain with temporal resolution $\Delta_{S}$ to the target domain with temporal resolution $\Delta_{T}$, where $\rho=\frac{\Delta_{T}}{\Delta_{S}}$.
Note that all the methods presented below have been derived in detail in the form of propositions in Section LABEL:sec:propositions.
Integral approximation adaptation method is derived from the Integral approximation method for obtaining discrete-time SSM from a continuous-time system model. Following Proposition LABEL:Integral_Method_Prop in Section LABEL:sec:propositions, the proposed mapping is
$\!\bm{H}_{v,(T)}$ $=\left(\bm{H}_{v,(S)}\right)^{\rho}$ (14a)
$\bm{H}_{k,(T)}$ $=(\bm{H}_{v,(T)}-\bm{I})(\bm{H}_{v,(S)}-\bm{I})^{-1}\bm{H}_{k,(S)}$ (14b)
for $k=s,f,r$.
Euler adaptation method is derived from the Euler approximation method for obtaining discrete-time SSM from a continuous-time system model. Following Proposition LABEL:Euler_Method_Prop in Section LABEL:sec:propositions, the proposed mapping is
$\!\bm{H}_{v,(T)}$ $=\bm{I}+\rho(\bm{H}_{v,(S)}-\bm{I})$ (15a)
$\bm{H}_{k,(T)}$ $=\rho\bm{H}_{k,(S)}$ (15b)
for $k=s,f,r$.
Expectation adaptation method is based on the assumption the expected values of the state variable vector between different time resolutions are the same. This adaptation method assumes that the ratio between the time steps ($\Delta$) of the lower and higher resolutions is an integer. Specifically, if $\rho>1,\rho\in\mathbb{Z}_{+}$, while if $\rho<1,1/\rho\in\mathbb{Z}_{+}$. Following Proposition LABEL:Method_Expectation_Prop in Section LABEL:sec:propositions the proposed mapping is
$\noindent\!\bm{H}_{v,(T)}$ $=\left(\bm{H}_{v,(S)}\right)^{\rho}$ (16a)
$\!\bm{H}_{k,(T)}$ =$=\begin{cases}(\sum^{\rho}_{j=1}\left(\bm{H}_{v,(S)}\right)^{\rho-j})\bm{H}_{k,(S)}&\rho\geq 1\\ (\sum^{1/\rho}_{j=1}\left(\bm{H}_{v,(T)}\right)^{1/\rho-j})^{-1}\bm{H}_{k,(S)}&\rho<1\end{cases}$ (16b)
for $k=s,f,r$.
Remark 1.
In the Integral and Expectation adaptation methods, $\left(\bm{H}_{v,(S)}\right)^{\rho}$ represents a matrix raised to a possibly fractional power which can be computed using the Schur–Padé algorithm [MatrixFractionalPower]. In general, fractional powers of a real-valued matrix may be complex-valued. By restricting the parameters of the adLIF neuron (1) to the ranges $\alpha,\beta,a\in(0,1)$, the possibly fractional power $\rho=1/\bar{\rho}$ of the real matrix $\bm{H}_{v,(S)}$ used in our experiments remains real-valued [MatrixFractionalPower].
4.4 A Simple Benchmark Method: Time-constant Adaptation Method
We now introduce a simple method that will serve as a benchmark for our proposed methods. This benchmark method is motivated by the explicit dependence of various neuron parameters on the sampling period in the popular LIF neuron and adaptive LIF neuron models, and has been utilized in the literature for the LIF neuron [he2020comparing]. Assume that the elements of the matrices $\bm{H}_{k,(S)}$ for $k=v,s,f,r$ belong to one of three groups:
  1. i
    in the form of $h_{(S)}=\exp^{-\Delta_{S}/\tau}$ where $\tau\in\mathbb{R}_{+}$;
  2. ii
    linear transformation of $h_{(S)}$;
  3. iii
    has no explicit dependency on $\Delta_{S}$.
The time-constant adaptation methods obtains $\bm{H}_{k,(T)}$ for $k=v,s,f,r$, by replacing elements of $\bm{H}_{k,(S)}$ belonging to group i, ii, iii by:
  1. i
    $h_{(T)}=\left(h_{(S)}\right)^{\rho}$
  2. ii
    the same linear transformation of $h_{(T)}$
  3. iii
    rest of the terms are unchanged.
To illustrate the method, we now look into the popular adaptive LIF neuron from Section 3.1. In the adaptive LIF neuron, all of the three types of elements are present. Here, $\alpha_{(T)}=\alpha_{(S)}^{\rho}$ and $\beta_{(T)}=\beta_{(S)}^{\rho}$ due to their explicit exponential dependence on the step length $T$. The other parameters are either scaled accordingly due to the dependence with the already scaled parameters, for example $-(1-\alpha_{(T)})$ or kept the same (in the case of the parameters $a$ and $b$).